A FAMILY INVEST SIX THOUSAND DOLLARS IN AN ACCOUNT PAYING THREE PERCENT COMPOUNDED MONTHLY. HOW MUCH IS IN THE ACCOUNT AFTER FOUR MONTHS?
Answer by mananth(16946)

Find Amount
Principal P = 6000
Amount= A
years=n 0.33
compounded 12 times a year t
Rate = 4.00 0.04
Amount = P*((n+r)/n)^n*t

Amount = = 6000 *( 1 + 0.04 /t)^ 0.33 * 12
Amount = 6000 *( 1 + 0 )^ 4
6000 *( 1 )^ 4
Amount = 6080.40